OpenStudy (anonymous):

Please help! Perform indicated operation and leave answer in a+bi form :/ (7+2i)/(4-3i) a. 22/25-13/25i b. 34/25-13/25i c. 34/25+29/25i d. 22/25+29/25i

OpenStudy (unklerhaukus):

\[\frac{7+2i}{4-3i} =\]\[=\frac{7+2i}{4-3i}\times1 \]\[=\frac{7+2i}{4-3i}\times\frac{4+3i}{4+3i} \]\[=\frac{(7+2i)(4+3i)}{4^2-3^2i^2}\]\[=\]

OpenStudy (anonymous):

d. Thanks! :)